Highway Express has paid annual dividends of $1.32, $1.33, $1.38, $1.40, and $1.42 over the past five years, respectively. What is the average dividend growth rate?

=1.85%

Please check the answers and show all work typed out. No excel or grid style please as I am on mobile.
Answer #1

we have to first compute annual divided growth rate

Year 2 dividend growth = (1.33-1.32)/1.32 = 0.76%

Year 3 dividend growth = (1.38-1.33)/1.33= 3.76%

Year 4 dividend growth = (1.4-1.38)/1.38=1.45%

Year 5 dividend growth = (1.42-1.4)/1.4 = 1.43%

So 4 year total growth = 0.76%+3.76%+1.45%+1.43%=7.39%

Average growth rate = 7.39%/4 = .85%